Assume that there is a 7% rate of disk drive failure in a year a all your computer data is stored on a hard disk drive with a copy stored on a second hard disk dtrive, what is the probablay that duning a year you can avoid catastrophe with at s on workling drive b. t copies of all your computer data are stored on three independent hard disk drives, whest is the probabilty hat during a year you can avoid catastrophe with at least one working drive? (Round to four decimal places as needed.) b. With three hard disk drives, the probablily that catastrophe can be avoided is (Round to sx decimal places as needed.) 3 caps lock
Answer #1

a)P(at least one disk works) =1-P(noen of 2 hard drive works) =1-(0.07)2 =0.9951

b)P(at least one disk works) =1-P(noen of 3 hard drive works) =1-(0.07)3 =0.999657